At the Dobren company, an all-day study revealed the sources of noise following: 0.5 h of 98 dBA; 1h less than 90 dBA; 3.5 h of 92 dBA; 3 h of 102 dBA Does the company comply with current regulations? What is the exposure dose? What is the TWA noise level?
Answer #1

In the United States, the OSHA (Occupational Safety and Health Administration) provides regulations for limiting noise exposure to workers. With difference in worker’s tasks and equipment in use, each industry has a different noise profile. However, noise becomes hazardous on exceeding 85 dBA for an eight-hour shift. Although, some industries have set the bar with a threshold of 90dBA. On exceeding 90 dBA, workers need to be provided with hearing protection. OSHA stipulates a permissible noise level for an eight-hour shift as 90 dBA, which is equivalent to an exposure dose of 100%. Considering a 5 dB exchange rate, the dose can be calculated as:

Dose = 100*(C1/T1 + C2/T2 + C3/T3 + Cn/Tn)

Where, Cn = Time spent each noise level; Tn = 8 /2 (L-90)/5

To calculate D in given case, we first need to calculate Tn as per given data.

T1 = 8/2(98-90)/5 = 2.63

T2 = 8/2(90-90)/5 = 8

T3 = 8/2(92-90)/5 = 6.06

T4 = 8/2(102-90)/5 = 1.51

Dose, D = 100*(0.5/2.63 + 1/8 + 3.5/6.06+3/1.51) = 100*(0.19+0.125+0.57+1.98) = 286.5%

Thus, Exposure Dose is 286.5%

Now, TWA = 16.61* Log10(D/100)+90

                   = 16.61* Log10 (286.5/100)+90

                   = 97.59

TWA noise level (Time Weighted Average) = 97.59